# The Discovery of Little Red Dots in the Local Universe: Signatures of Cool Gas Envelopes

### v1

@@ Abstract
JWST observations have revealed a population of high-redshift “little red dots” (LRDs) that challenge conventional AGN models. We report the discovery of three local LRDs at –, initially selected from the SDSS database, with follow-up optical/near-IR spectroscopy and photometry. They exhibit properties fully consistent with those of high-redshift LRDs, including broad hydrogen and helium emission lines, compact morphologies, V-shaped UV-optical SED, declining near-IR continua, and no significant variability. Two sources were targeted but not detected in X-rays with statistical significance. All three sources show blue-shifted He I absorption, while two exhibit H and Na D absorption lines. We detect full Balmer and Paschen line series in all three objects, along with abundant narrow [Fe II] emission in two. The emission line analyses suggest narrow lines originate from AGN-powered, metal-poor regions with minimal dust; broad lines come from inner regions with exceptionally high density or atypical dust properties; and [Fe II] emission arises from dense gas between broad and narrow-line regions. One of our objects, J1025+1402 (nicknamed The Egg), shows extremely high equivalent width Na D, K I, Fe II, and Ca II triplet absorption lines, suggesting the presence of a cool (5000 K), metal-enriched gas envelope. The optical/near-IR continua of these LRDs are also consistent with theoretical models featuring an atmosphere around black holes. The WISE-detected IR emission is consistent with weak dust emission of K. We propose a conceptual model consisting of a largely thermalized cool-gas envelope surrounding the central black hole and an extended emission line region with high-density outflowing gas to explain the observed properties of these local LRDs.
@@ 1 Introduction
The so-called “little red dots” (LRDs) have been among the most significant discoveries of the early years of JWST operations (e.g., I. Labbe et al., 2025; J. Matthee et al., 2024; X. Lin et al., 2024; H. B. Akins et al., 2024). These sources appear compact even in JWST imaging, and the majority of them (over 70–80%) exhibit broad Balmer emission lines with FWHM km s-1, indicative of black holes (BHs) of – (e.g., J. E. Greene et al., 2024; D. D. Kocevski et al., 2024; J. Zhang et al., 2025). They display V-shaped spectral energy distributions (SEDs), characterized by a blue UV continuum and a red optical continuum (D. D. Kocevski et al., 2024; D. J. Setton et al., 2024; R. E. Hviding et al., 2025). Moreover, a large fraction of these V-shaped objects present Balmer absorption on top of the broad emission line, which is rarely seen in other type-1 AGNs (e.g., J. Matthee et al., 2024; X. Lin et al., 2024; B. Wang et al., 2025a; F. D’Eugenio et al., 2025).
@@
Debates about the nature of LRDs are ongoing, particularly concerning the origin of their V-shaped SEDs, the spectral shape inflection near the Balmer break, and the prevalence of Balmer absorption (e.g., D. J. Setton et al., 2024; X. Lin et al., 2024). Early studies have shown that if the Balmer break arises from massive host galaxies, it would challenge our current understanding of structure formation (e.g., B. Wang et al., 2024; I. Labbe et al., 2024). Alternatively, if these features are attributed to reddened AGNs, it is difficult to account for the lack of strong rest-frame IR detections for most high-redshift LRDs, by JWST/MIRI, Spitzer, or ALMA (e.g., P. G. Pérez-González et al., 2024; C. C. Williams et al., 2024; Y. Ma et al., 2024; D. J. Setton et al., 2025). New models have started to explore the distribution of dust (Z. Li et al., 2025; K. Chen et al., 2025), along with new constraints on the dust content in LRDs from multi-wavelength data (C. M. Casey et al., 2024, 2025).
@@
The unusual continuum shape of LRDs has motivated multiple theoretical models (K. Inayoshi et al., 2025; D. Kido et al., 2025). Recent studies have proposed that BHs embedded in dense gas with a high covering fraction can explain the presence of both the observed Balmer breaks and absorption features (K. Inayoshi & R. Maiolino, 2024). Such models have been applied to several high-redshift LRDs and have achieved reasonably good fits to their spectra (X. Ji et al., 2025; A. J. Taylor et al., 2025a; A. de Graaff et al., 2025; R. P. Naidu et al., 2025). However, for sources with rest-frame IR detections, these models struggle to simultaneously reproduce the weak IR photometry and the shape of Balmer breaks. They typically require either an extremely high column density of hydrogen gas (e.g., ), exotic modified dust attenuation laws, or intrinsically red AGN SEDs (A. de Graaff et al., 2025). Even with these additional free parameters, challenges still remain for fully explaining the IR broadband photometry (e.g., A. de Graaff et al., 2025; R. P. Naidu et al., 2025; A. J. Taylor et al., 2025a).
@@
Currently, only a limited number of high-redshift LRDs have been detected with JWST/MIRI, while the vast majority remain either not covered by or undetected in the mid-IR (e.g., G. C. K. Leung et al., 2024; D. J. Setton et al., 2025; X. Lin et al., 2025; A. J. Taylor et al., 2025a). Most JWST/NIRSpec spectroscopy of LRDs at has been conducted in the prism mode, and high-resolution grating observations of several objects reveal only their brightest features. In contrast, LRDs at low redshifts can be observed with a variety of facilities/instruments at multiple wavelengths, with high spectral and spatial resolution at relatively low cost, providing deep insights into the nature of LRDs as a population. Rest-frame IR and high-resolution grating spectra have been obtained for a small number of luminous LRDs at cosmic noon (e.g., I. Juodžbalis et al., 2024; B. Wang et al., 2025a; A. de Graaff et al., 2025). Systematic searches for LRDs at using wide-field sky surveys are now underway and have already demonstrated successful spectroscopic follow-up with ground-based facilities (R. Lin et al., 2025; Y. Ma et al., 2025; Euclid Collaboration et al., 2025). However, photometrically selected LRDs at cosmic noon are still relatively faint ( mag), making follow-up observations challenging with current ground-based facilities.
@@
Identifying LRDs in the local Universe opens a valuable window into understanding the nature of this population. Recent observations and theoretical predictions suggest that their number density declines significantly at , dropping by 1–2 dex from to (D. D. Kocevski et al., 2024; Y. Ma et al., 2025; K. Inayoshi, 2025). This implies that local LRDs are rare and are linked to physical conditions that may be more common in the early Universe. Their discovery at low redshift would offer critical insights into the evolutionary pathways of LRDs and their connection to BH growth across cosmic time. Moreover, local LRDs may enable spatially resolved observations of themselves, their host galaxies, and their environments, revealing their structure, kinematics, and dust/gas content. These observations would not only be cost-effective but also may unravel details that remain challenging to obtain from their high-redshift counterparts, even with JWST.
@@
In this paper, we present the search and discovery of three local LRDs at =0.1–0.2 from the SDSS spectroscopic database, followed by high signal-to-noise radio ground-based spectroscopic observations. The paper is organized as follows. In §2 we introduce the selection criteria we used to search the SDSS spectral database for local LRDs. We present the SDSS data and follow-up observations on the three selected local LRDs in §3. In §4 we describe their overall properties and spectral features, which are fully consistent with those of JWST-discovered high-redshift LRDs. In §5, we analyze their emission line properties and infer the physical conditions. We then conduct a detailed case study of one object in §6. Finally, in §7 we interpret the observed features to understand their physical nature. Throughout this work, we adopt the AB magnitude system for all photometric measurements. All equivalent widths (EWs) are reported in the rest frame. A flat CDM cosmology is assumed, with , , and .
@@ 2 Selection and sample from SDSS
In the literature, a wide range of selection criteria have been used to define the LRD population (e.g., K. N. Hainline et al., 2025; R. E. Hviding et al., 2025). In this paper, we adopt a narrow definition of LRDs, referring specifically to objects that simultaneously satisfy the following criteria: (1) presence of broad Balmer emission lines; (2) compact morphology in the rest-frame optical, particularly in photometric bands dominated by H emission; and (3) a V-shaped SED, characterized by a blue UV continuum slope and red optical continuum slope. A detailed description of the selection of local LRDs will be presented in an upcoming paper. Here, we provide a summary of the procedure.
@@ 2.1 Selection
We first built a library of high-redshift LRDs with well-characterized UV-to-optical spectra. We adopted the compilation of high-redshift LRDs from D. J. Setton et al. (2024), which exhibit clear V-shaped SEDs with inflection points near the Balmer limit. Their JWST/NIRSpec PRISM spectra were obtained from the DAWN JWST Archive 111https://dawn-cph.github.io/dja/.
@@
We then searched for local LRD candidates in the SDSS DR17222https://www.sdss4.org/dr17/ spectroscopic database (K. N. Abazajian et al., 2009; Abdurro’uf et al., 2022). We started with the GUVmatch333https://archive.stsci.edu/hlsp/guvcat catalog, which cross-matches GALEX far-UV (FUV) and near-UV photometry (NUV) with SDSS DR14 ugriz photometry. We shifted the high-redshift LRD PRISM spectra to –0.5 to serve as templates for our local LRD search. This redshift range was chosen to ensure that the H emission line falls within the SDSS spectroscopic wavelength coverage. We then re-projected the template spectra to GALEX and SDSS photometry to obtain the expected broadband photometry of local LRDs. Using these synthetic colors, we searched for candidates in GUVmatch with similar SED shapes. Specifically, we computed the Euclidean distance in magnitude space between the observed photometry and each individual template. This distance was defined as the root mean square of the magnitude differences across the GALEX and SDSS bands, normalized by the -band flux. Then, we selected sources with a distance of less than one magnitude from any of the templates. Finally, we imposed a compactness criterion by requiring the SDSS-reported Petrosian radii (V. Petrosian, 1976) in both the g and r bands to be smaller than 1.8 arcsec (physical scale of 3.3 kpc at ). This morphological criterion was relatively permissive, and variations in the threshold do not affect the selection results.
@@
After the photometric selection, we retrieved the SDSS spectra of the selected candidates. We examined their emission line properties using the SPZLINE extension, which provides automated emission-line fits from the SDSS pipeline (A. S. Bolton et al., 2012). We first required the EWs of [O III] 5008 to exceed 10 Å, and the line ratio [O III] 5008/[O II] 3727 (O32) to be greater than 10. The first criterion effectively excludes stellar objects, quiescent galaxies, and post-starburst galaxies, which could exhibit strong Balmer breaks and are prevalent in the local Universe. The second criterion ensures that the selected galaxies have physical conditions similar to those of high-redshift galaxies.
@@
To further remove contaminants and confirm the presence of red optical continua, we inspected the continuum underlying emission lines in the SPZLINE extension. We required the continuum under H to be redder than that under [O III] 5008, i.e., in space. We also required the [O III] 5008 continuum to be redder than that under [O II] 3727.
Although we define LRDs by the presence of broad H lines, up to this stage, the broad-line criterion has not yet been explicitly applied during the selection process above. Nonetheless, all targets selected based on the above criteria have already exhibited clear broad H emission in their SDSS spectra. We therefore obtained a sample of local sources that meet all the criteria defining LRDs from the SDSS DR17 spectroscopic database.
@@ 2.2 Sample from SDSS
In fact, the local LRDs presented in this paper were originally selected as quasar candidates (G. T. Richards et al., 2002). The SDSS target selection flag for J1025+1402 and J1022+0841 is QSO_HIZ, as their colors meet the criteria for selecting high-redshift quasars, primarily at . Both objects show rising flux toward and peaking in the band, followed by a decline toward the bands, resembling the spectral shape of a Lyman break with a blue UV continuum, characteristic of quasars at . J1047+0739 is flagged as both QSO_CAP and QSO_HIZ: the former because its color meets the selection criteria for quasars at , and the latter because it satisfies a relaxed criterion for high-redshift quasars. These flags indicate that the three selected objects are outliers in both the and color spaces, making them easily misclassified during color selection.
Our LRD selection is incomplete, since it requires that the object has an existing SDSS spectrum. The SDSS main galaxy survey is biased against LRDs because of their compact morphology, which is one of the reasons that they are labeled as quasar candidates. Therefore, we estimate a conservative lower limit on the number density of local LRDs at =0–0.5, based on the effective SDSS survey area of 9,376 deg2. The discovery of these three objects yields an estimated lower limit of approximately within this redshift range. Our estimated number density at =0–0.5 is more than several orders of magnitude higher than the extrapolation of the log-normal distribution fitted to the LRD population at 4–7 (K. Inayoshi, 2025). However, this large discrepancy arises in the exponentially damped tail of the assumed distribution, where such divergence is expected due to the steep functional form and limited constraints at low redshifts. We find that a modest increase in the dispersion (by 10–20%) of the log-normal model from K. Inayoshi (2025) yields a redshift evolution of the LRD abundance consistent not only with measurements at 2–3 (Y. Ma et al., 2025), but also with our observations at =0–0.5. Nonetheless, this trend suggests that the BH accretion mode giving rise to LRDs, while rare in the local Universe, is not exclusively unique to the early Universe.
@@
From the SDSS spectra, J1025+1402 shows clear blueshifted H absorption, while J1022+0841 exhibits tentative but unresolved H absorption. Both features are confirmed by our spectroscopic follow-up with higher spectral resolution (see §4.3). J1025+1402 and J1047+0739 were identified by Y. I. Izotov & T. X. Thuan (2008) as metal-poor AGNs in SDSS DR5, with 12 + (O/H) = 7.36 and 7.99, respectively. C. Simmonds et al. (2016) further noted that they are weak in X-rays, revealed by the non-detection of Chandra. C. J. Burke et al. (2021) found no significant variability in either optical broadband photometry or broad H luminosity.

### v2

@@ Abstract
JWST observations have revealed a population of high-redshift “little red dots” (LRDs) that challenge conventional AGN models. We report the discovery of three local LRDs at –, initially selected from the SDSS database, with follow-up optical/near-IR spectroscopy and photometry. They exhibit properties fully consistent with those of high-redshift LRDs, including broad hydrogen and helium emission lines, compact morphologies, V-shaped UV-optical SED, declining near-IR continua, and no significant variability. Two sources were targeted but not detected in X-rays with statistical significance. All three sources show blue-shifted He I absorption, while two exhibit H and Na D absorption lines. We detect full Balmer and Paschen line series in all three objects, along with abundant narrow [Fe II] emission in two. The emission line analyses suggest narrow lines originate from AGN-powered, metal-poor regions with minimal dust; broad lines come from inner regions with exceptionally high density or atypical dust properties; and [Fe II] emission arises from dense gas between broad and narrow-line regions. One of our objects, J1025+1402 (nicknamed The Egg), shows extremely high equivalent width Na D, K I, and Ca II triplet absorption lines, along with other potential low-ionization absorption features, suggesting the presence of a cool (5000 K), metal-enriched gas envelope. The optical/near-IR continua of these LRDs are also consistent with theoretical models featuring an atmosphere around black holes. The WISE-detected IR emission is consistent with weak dust emission of K. We propose a conceptual model consisting of a largely thermalized cool-gas envelope surrounding the central black hole and an extended emission line region with high-density outflowing gas to explain the observed properties of these local LRDs.
@@ 1 Introduction
The so-called “little red dots” (LRDs) have been among the most significant discoveries of the early years of JWST operations (e.g., Labbe2025; Matthee2024; Lin2024; Akins2024). These sources appear compact even in JWST imaging, and the majority of them (over 70–80%) exhibit broad Balmer emission lines with FWHM km s-1, indicative of black holes (BHs) of – (e.g., Greene2024; Kocevski2024; Zhang2025). They display V-shaped spectral energy distributions (SEDs), characterized by a blue UV continuum and a red optical continuum (Kocevski2024; Setton2024; Hviding2025). Moreover, a large fraction of these V-shaped objects presents Balmer absorption on top of the broad emission line, which is rarely seen in other type-1 AGNs (e.g., Matthee2024; Lin2024; Wang2025; DEugenio2025).
@@
Debates about the nature of LRDs are ongoing, particularly concerning the origin of their V-shaped SEDs, the spectral shape inflection near the Balmer break, and the prevalence of Balmer absorption (e.g., Setton2024; Lin2024). Early studies have shown that if the Balmer break arises from massive host galaxies, it would challenge our current understanding of structure formation (e.g., Wang2024; Labbe2024). Alternatively, if these features are attributed to reddened AGNs, it is difficult to account for the lack of strong rest-frame IR detections for most high-redshift LRDs, by JWST/MIRI, Spitzer, or ALMA (e.g., Perez-Gonzalez2024; Williams2024; Ma2024; Xiao2025; Setton2025). New models have started to explore the distribution of dust (Li2025; Chen2025), along with new constraints on the dust content in LRDs from multi-wavelength data (Casey2024; Casey2025). In addition to their weak infrared and sub-millimeter emission, the weak X-ray (Yue2024; Ananna2024; Sacchi2025) and radio emission (Mazzolari2024; Perger2025), further set LRDs apart from typical type-1 AGNs (Kokubo2024; Maiolino2025).
@@
The unusual continuum shape of LRDs has motivated multiple theoretical models (Inayoshi2025b; Kido2025). Recent studies have proposed that BHs embedded in dense gas with a high covering fraction can explain the presence of both the observed Balmer breaks and absorption features (Inayoshi2024). Such models have been applied to several high-redshift LRDs and have achieved reasonably good fits to their spectra (Ji2025; Taylor2025; deGraaff2025; Naidu2025). However, for sources with rest-frame IR detections, these models struggle to simultaneously reproduce the weak IR photometry and the shape of Balmer breaks. They typically require either an extremely high column density of hydrogen gas (e.g., ), exotic modified dust attenuation laws, or intrinsically red AGN SEDs (deGraaff2025). Even with these additional free parameters, challenges remain for fully explaining the IR broadband photometry (e.g., deGraaff2025; Naidu2025; Taylor2025).
@@
Currently, only a limited number of high-redshift LRDs have been detected with JWST/MIRI, while the vast majority remain either not covered by or undetected in the mid-IR (e.g., Leung2024; Setton2025; Lin2025; Taylor2025). Most JWST/NIRSpec spectroscopy of LRDs at has been conducted in the prism mode, and high- or medium-resolution grating observations of several objects reveal only their brightest features. In contrast, LRDs at low redshifts can be observed with a variety of facilities/instruments at multiple wavelengths, with high spectral and spatial resolution at relatively low cost, providing deep insights into the nature of LRDs as a population. Rest-frame IR and high-resolution grating spectra have been obtained for a small number of luminous LRDs at cosmic noon (e.g., Juodzbalis2024; Wang2025; deGraaff2025). Systematic searches for LRDs at using wide-field sky surveys are now underway and have already demonstrated successful spectroscopic follow-up with ground-based facilities (RLin2025; Ma2025; LRD_EuclidCollaboration2025). However, photometrically selected LRDs at cosmic noon are still relatively faint ( mag), making follow-up observations challenging with current ground-based facilities.
@@
Identifying LRDs in the local Universe opens a valuable window into understanding the nature of this population. Although the measurements of the evolution of LRD number density across cosmic time vary across different photometric and spectroscopic selections, recent observations and models show a decline in the LRD number density toward low redshift. (Ma2025; LRD_EuclidCollaboration2025; Loiacono2025; Inayoshi2025a). This implies that local LRDs are rare and are linked to physical conditions that may be more common in the early Universe. Their discovery at low redshift would offer critical insights into the evolutionary pathways of LRDs and their connection to BH growth across cosmic time. Moreover, local LRDs may enable spatially resolved observations of themselves, their host galaxies, and their environments, revealing their structure, kinematics, and dust/gas content. These observations would not only be cost-effective but also may unravel details that remain challenging to obtain from their high-redshift counterparts, even with JWST.
@@
In this paper, we present the search and discovery of three local LRDs at =0.1–0.2 from the SDSS spectroscopic database, followed by high signal-to-noise (S/N) ratio ground-based spectroscopic observations. The paper is organized as follows. In §2 we introduce the selection criteria we used to search the SDSS spectral database for local LRDs. We present the SDSS data and follow-up observations on the three selected local LRDs in §3. In §4 we describe their overall properties and spectral features, which are fully consistent with those of JWST-discovered high-redshift LRDs. In §5, we analyze their emission line properties and infer the physical conditions. We then conduct a detailed case study of one object in §6. Finally, in §7 we interpret the observed features to understand their physical nature. Throughout this work, we adopt the AB magnitude system for all photometric measurements. All wavelengths and line names are given in the vacuum frame. All equivalent widths (EWs) are reported in the rest frame. A flat CDM cosmology is assumed, with , , and .
@@ 2 Selection and sample from SDSS
In the literature, a wide range of selection criteria have been used to define the LRD population (e.g., Hainline2025; Hviding2025). In this paper, we adopt a narrow definition of LRDs, referring specifically to objects that simultaneously satisfy the following criteria: (1) presence of broad Balmer emission lines; (2) compact morphology in the rest-frame optical, particularly in photometric bands dominated by H emission; and (3) a V-shaped SED, characterized by a blue UV continuum slope and red optical continuum slope. A detailed description of the selection of local LRDs will be presented in an upcoming paper. Here, we provide a summary of the procedure.
@@ 2.1 Selection
We first built a library of high-redshift LRDs with well-characterized UV-to-optical spectra. We adopted the compilation of high-redshift LRDs from Setton2024, which exhibit clear V-shaped SEDs with inflection points near the Balmer limit. Their JWST/NIRSpec PRISM spectra were obtained from the DAWN JWST Archive 111https://dawn-cph.github.io/dja/.
@@
We then searched for local LRD candidates in the SDSS DR17222https://www.sdss4.org/dr17/ spectroscopic database (Abazajian2009; Abdurrouf2022). We started with the GUVmatch catalog (GUVcat_DOI), which cross-matches GALEX far-UV (FUV) and near-UV photometry (NUV) with SDSS DR14 ugriz photometry. We shifted the high-redshift LRD PRISM spectra to –0.5, serving as templates for our local LRD search. This redshift range was chosen to ensure that the H emission line falls within the SDSS spectroscopic wavelength coverage. We then re-projected the template spectra to GALEX and SDSS photometry to obtain the expected broadband photometry of local LRDs. Using these synthetic colors, we searched for candidates in GUVmatch with similar SED shapes. Specifically, we computed the Euclidean distance in magnitude space between the observed photometry and each individual template. This distance was defined as the root mean square of the magnitude differences across the GALEX and SDSS bands, normalized by the -band flux. Then, we selected sources with a distance of less than one magnitude from any of the templates. Finally, we imposed a compactness criterion by requiring the SDSS-reported Petrosian radii (Petrosian1976) in both the g and r bands to be smaller than 1.8 arcsec (physical scale of 3.3 kpc at ). This morphological criterion was relatively permissive, and variations in the threshold do not affect the selection results.
@@
After the photometric selection, we retrieved the SDSS spectra of the selected candidates. We examined their emission line properties using the SPZLINE extension, which provides automated emission-line fits from the SDSS pipeline (Bolton2012). We first required the EWs of [O III] 5008 to exceed 10 Å, and the line ratio [O III] 5008/[O II] 3730 (O32) to be greater than 10. The first criterion effectively excludes stellar objects, quiescent galaxies, and post-starburst galaxies, which could exhibit strong Balmer breaks and are prevalent in the local Universe. The second criterion ensures that the selected galaxies have physical conditions similar to those of high-redshift galaxies (e.g., Shapley2023; Sanders2023; Shapley2025).
@@
To further remove contaminants and confirm the presence of red optical continua, we inspected the continuum underlying emission lines in the SPZLINE extension. We required the continuum under H to be redder than that under [O III] 5008, i.e., in space. We required the [O III] 5008 continuum to be redder than that under [O II] 3730.
After obtaining the initial parent sample based on the automated pipeline products, we visually inspected the candidates to exclude objects affected by pipeline uncertainties. For example, we removed cases where the more precise re-measured [O II] flux remained clearly stronger than [O III] , or where emission lines had low S/N and were dominated by noise. This process resulted in a parent sample of 40 objects.
Although the broad-line H criterion, which is one of the defining characteristics of LRDs, was not explicitly applied up to this stage, all selected targets already show clear broad H emission in their SDSS spectra. We note, however, that not all objects in the parent broad-line sample satisfy the LRD definition. Many sources in the parent sample show very strong [N II] or Mg II emission accompanied by relatively red optical continua, in which case the spectral shapes are more naturally explained by type-1 AGNs with evolved stellar populations. In the end, we select three objects as the most representative LRDs.
@@ 2.2 Sample from SDSS
The local LRDs presented in this paper were originally selected as quasar candidates (Richards2002). The SDSS target selection flag for J1025+1402 and J1022+0841 is QSO_HIZ, as their colors meet the criteria for selecting high-redshift quasars, primarily at . Both objects show rising flux toward and peaking in the band, followed by a decline toward the bands, resembling the spectral shape of a Lyman break with a blue UV continuum, characteristic of quasars at . J1047+0739 is flagged as both QSO_CAP and QSO_HIZ: the former because its color meets the selection criteria for quasars at , and the latter because it satisfies a relaxed criterion for high-redshift quasars. The misidentification indicates that the three selected objects are outliers in both the and color spaces, making them easily misclassified during color selection.
@@
From the SDSS spectra, J1025+1402 shows clear blueshifted H absorption, while J1022+0841 exhibits tentative but unresolved H absorption. Both features are confirmed by our spectroscopic follow-up with higher spectral resolution (see §4.3 and Figure 5). J1025+1402 and J1047+0739 were identified by Izotov2008 as metal-poor AGN in SDSS DR5, with 12 + (O/H) = 7.36 and 7.99, respectively. Simmonds2016 further noted that they are weak in X-rays, revealed by the non-detection of Chandra. Burke2021 found no significant variability in either optical broadband photometry or broad H luminosity.
2.3 Number density of LRDs at
Our LRD selection at is incomplete. First of all, it requires that the object has an existing SDSS spectrum. The SDSS main galaxy survey is biased against LRDs because of their compact morphology, making them less likely to be prioritized for spectroscopic follow-up as galaxies. Their compactness is also one of the reasons that they are labeled as quasar candidates. In addition, although our [O III] EW criteria effectively exclude stars, quiescent, and post-starburst galaxies at low redshifts, they may also omit potential LRD candidates with intrinsically weak [O III] emission. Indeed, several high-redshift LRDs exhibit low [O III] EWs, such as Abell2744-QSO1 (4.5 Å; Ji2025), MoM-BH∗-1 (3 Å; Naidu2025), and The Cliff ( Å; deGraaff2025).
Therefore, we estimate a conservative lower limit on the number density of local LRDs at =0–0.5, based on the effective SDSS survey area of 9,376 deg2. The discovery of these three objects yields an estimated lower limit of approximately within this redshift range. Our estimated number density at =0–0.5 is more than several orders of magnitude higher than the extrapolation of the log-normal distribution fitted to the LRD population at 4–7 (Inayoshi2025a). However, this large discrepancy arises in the exponentially damped tail of the assumed distribution, where such divergence is expected due to the steep functional form and limited constraints at low redshifts. We find that a modest increase in the dispersion (by 10–20%) of the log-normal model from Inayoshi2025a yields a redshift evolution of the LRD abundance consistent not only with measurements at 2–3 (Ma2025), but also with our observations at =0–0.5.
The evolution of LRD number density across cosmic time remains uncertain, with different studies reporting varying trends. Photometric selection and spectroscopic follow-ups by Ma2025 indicate a substantial decline at , with the number density decreasing by 1–2 dex from to . In contrast, photometric selection in LRD_EuclidCollaboration2025 suggests a milder evolution, with LRD abundance rising to before declining at lower redshift. Based on the discovery of three LRDs at , Loiacono2025 implies that LRD number density remains significant at cosmic noon, only 2–3 times lower than that of UV-selected quasars. Nevertheless, both Ma2025 and LRD_EuclidCollaboration2025 indicate a decreasing number density toward the local Universe. This trend, together with the discovery of the three local LRDs presented in this work, suggests that the BH accretion mode giving rise to LRDs, while rare in the local Universe, is not exclusively unique to the early Universe.